• Is Industrial Economics different from Industrial Organisation? Justify your answer.

Below is a list of three (3) different markets. The market share of the firms in each market is given. Market A: 40%; 20%; 23%; 9%; 8%.

Market B: 26%; 24%; 23%; 27%

.Market C: 80%; 10%; 6%; 3%;1%. •

Calculate the 3-firm Concentration Ratio, CR(3) of each market. Interpret your results. •

Calculate the Herfindahl-Hirschman Index (HHI) of each market. Interpret your results. •

Which one of the two methods of market concentration above is more robust (accurate)? Justify your answer.


1
Expert's answer
2021-06-14T13:14:09-0400

Solution:

Industrial Economics and Industrial Organization are similar in nature.

Industrial Economics is the study of firms, industries and markets and it studies firms of all sizes. It also provides insights into how firms organize their activities, including considering their motivation and assessing whether a market is competitive or not.

Industrial Organization is the economic field that deals with the study of firms strategic behaviors and their interaction to establish the structure of markets. It is concerned with the workings of markets and industries, especially the way firms compete with each other and how the markets operate.

 

A 3-firm Concentration Ratio is calculated by adding up the market share taken up by the three biggest firms:

Market A = "40\% + 20\% + 23\% = 83\%"

3-firm Concentration Ratio of Market A = "83\%"

This means that an oligopoly exists in the market since the top three firms in the market account for more than 60% of the total market sales.

 

Market B = "26\% + 24\% + 23\% = 73\%"

3-firm Concentration Ratio of Market B = "73\%"

This also means that an oligopoly exists in the market since the top three firms in the market account for more than 60% of the total market sales.

 

Market C = "80\% + 10\% +6\% = 96\%"

3-firm Concentration Ratio of Market B = "96\%"

This also means that an oligopoly exists in the market since the top three firms in the market account for more than 60% of the total market sales.

 

Calculating the Herfindahl-Hirschman Index (HHI) of each market, you square the market share of each firm competing in a market and then summing the resulting numbers.

Market A HHI = 402 + 202 + 232 + 92 + 82 = 1600 + 400 + 529 + 81 + 64 = 2,674

Since the HHI is greater than 2,500, then this means that the market is highly concentrated, hence experiences low competition.

 

Market B HHI = 262 + 242 + 232 + 272 = 676 + 576 + 529 + 729 = 2,510

Since the HHI is greater than 2,500, then this means that the market is highly concentrated, hence experiences low competition.


Market A HHI = 802 + 102 + 62 + 32 + 12 = 6,400 + 100 + 36 + 9 + 1 = 6,546

Since the HHI is greater than 2,500, then this means that the market is highly concentrated and hence experiences low competition.


The Herfindahl-Hirschman Index (HHI) is a more robust and accurate measurement method of calculating market concentration. This is because it is expressed as the sum of squared market shares of all firms in an industry, as firms are weighted according to their size. It is a more precise measure because it takes into account all companies in the industry